OpenStudy (anonymous):

find m<1 in rectangle RSTW. I drew it in the comments...PLEASE HELPPP :(

OpenStudy (goformit100):

Where is the pic

OpenStudy (anonymous):



OpenStudy (shubhamsrg):



OpenStudy (anonymous):

so u just add 17 + 17?

OpenStudy (shubhamsrg):

17+17 would be the angle adjacent to <1> , to find <1>, use the fact that sum of angles in a triangle = 180 degrees

OpenStudy (anonymous):

ok so do i do the sum interior angle theorem? or do i just add 180 to 17 + 17, like 180+17+17?

OpenStudy (shubhamsrg):

hmm <1> + 17 + 17 = 180, thats what I meant from here, find <1>

OpenStudy (anonymous):

<1> + 34 = 180 -34 -34 <1> = 146

OpenStudy (shubhamsrg):

bingo! ^_^

OpenStudy (anonymous):

thank you so much!!

OpenStudy (shubhamsrg):

glad to help.. hmm and math doesn't such that much! :|

OpenStudy (anonymous):

lol i guess i suck at math which is why i hate it

OpenStudy (shubhamsrg):

hmm..

OpenStudy (shubhamsrg):

no one is good at maths , everyone learns something new every other time :)

OpenStudy (anonymous):

plus my teacher gave us so much homework...so i was really pissed when i made an account on here

OpenStudy (shubhamsrg):

aha..nvm hope your hatred gets reduced someday! ^^

OpenStudy (anonymous):

yeah hopefully someday :D